A loan of $40,000 is made at 4.75% interest , compounded annually after how many years will the amount due reach 80,000
Answer by lwsshak3(11628)

A loan of $40,000 is made at 4.75% interest , compounded annually after how many years will the amount due reach 80,000
compound interest formula: A=P(1+r/n)^nt, P=initial investment, r=interest rate, n=number of compounding periods per year, A=amount after t-years.
..
For given problem:
P=40000
r=.0475
n=1
..
80000=40000(1+.0475/1)^1*t
80000/40000=2=(1.0475)^t
(1.0475)^t=2
take log of both sides
tlog(1.0475)=log(2)
t=log(2)/log(1.0475)
t≈14.9
after how many years will the amount due reach 80,000? approximately 15